Parth (parthkohli):

I have heard that \(\log x\) is actually supposed to be \(\ln x\). I don't think that's true.

Parth (parthkohli):

Because ln is different, and log base 10 is different.

OpenStudy (ash2326):

log to the base e is ln \[\log_e x =\ln x\]

Parth (parthkohli):

Yeah, I knew it :D

Parth (parthkohli):

Now I'd accuse him for confusing me ;)

OpenStudy (ash2326):

and log x is not necessarily 10, the base can be anything greater than 0 but not 1 \[\log_b a,\ b>0, b\ne 1\]

OpenStudy (anonymous):

lnx = 2.303log(10)x

Parth (parthkohli):

Yeah @UnkleRhaukus you meant \(\ln x = \log _e x\) :P

OpenStudy (unklerhaukus):

yeah i did,

Parth (parthkohli):

But log x is supposed to be \(\log _{10} x\), when the base is not given

OpenStudy (unklerhaukus):

for an unspecified base i would assume base 10

Parth (parthkohli):

Yep :)

OpenStudy (unklerhaukus):

unless when using Wolfram

Parth (parthkohli):

Wolfram gets confused lol

OpenStudy (ash2326):

It's the properties of log, you can use any base as long you don't have it specified. Most of the log tables are either to base 10 or e

Parth (parthkohli):

But an unspecified base is def. 10

OpenStudy (ash2326):

Nope, that's not necessary. Because base 10 is common, we use it mostly

OpenStudy (anonymous):

If base is not given, then it is assumed as Log having 10 base...